# On completeness of modified wave operators for defocusing NLS

### v1

@@ Abstract.
The solution to this problem can be reformulated in terms of the completeness of wave operators. For $n=1$, we obtain a complete answer, even for large initial data. For $n=2$ completeness is established under suitable control of the $L^{\infty}$ norm of the solution $u(t).$
We prove that the $L^{\infty}$ bound of the solution $u(t)$ matches the decay rate of the linear Schrödinger equation only in the case $n=1.$ For $n\geq 2,$ a similar bound holds only for the $L^{2n/(n-1)}$ norm of $u(t).$
In the case where no control on the $L^{\infty}$ norm is available, we establish similar asymptotic behavior by replacing the $L^{2}$ norm in (0.1) with smaller Besov-type spaces.
@@ 1.2. Purpose of the work
Our goal is to solve Q2) without smallness on the initial data at least for $n=1.$
@@
Let us mention that in the case $n=1$ the question Q1) is treated in the preprint [25].
@@ 2. Main results
Next, our goal is to prove some $L^{q}$ decay estimates closely connected with the key assumption (2.1) in Theorem 2.1.
@@
If $n=1$ and $u_{0}\in\mathcal{H}^{1,1},$ then the solution to (1.1) satisfies
(2.12) $\|u(t)\|_{L^{\infty}}\leq Ct^{-1/2},t\geq 1.$
For $n\geq 2$ we can control the $L^{q}$ norm of the solution, where $q$ is limited by an index associated with Caffarelli - Kohn - Nirenberg inequality.
@@
If $n\geq 2$, $u_{0}\in\mathcal{H}^{1,1},$ and $q\in[2,2n/(n-1)],$ then the solution to (1.1) satisfies
(2.13) $\|u(t)\|_{L^{q}}\leq Ct^{-n/2+n/q},t\geq 1.$
We have another estimate provided initial data are radial.
Theorem 2.5.
If $n\geq 2$ and $u_{0}\in\mathcal{H}^{1,1}_{rad},$ then the solution to (1.1) satisfies
(2.14) $\||x|^{n/2}u(t)\|_{L^{\infty}}\leq C,t\geq 1.$
Again we give a brief idea of the proofs of these $L^{q}$ bounds.
We use pseudoconformal transformation (2.10). Then the results of Theorems 2.3, 2.4, 2.2 follow from the following.
Theorem 2.6.
If $n=1$ and $u_{0}\in\mathcal{H}^{1,1},$ then the solution to (2.11) satisfies
(2.15) $\|v(t)\|_{L^{\infty}}\leq C,t\in[0,1).$
Theorem 2.7.
If $n\geq 2$, $u_{0}\in\mathcal{H}^{1,1}$ and $q\in[2,2n/(n-1)],$ then the solution to (2.11) satisfies
(2.16) $\|v(t)\|_{L^{q}}\leq C,t\in[0,1).$
Theorem 2.8.
If $n\geq 2$ and $u_{0}\in\mathcal{H}^{1,1}_{rad},$ then the solution to (2.11) satisfies
(2.17) $\||x|^{n/2}v(t)\|_{L^{\infty}}\leq C,t\in[0,1).$
Then the $L^{q}$ bounds are obtained by using the energy estimate for $v$ (see Lemma 6.2)
(2.18) $\int_{\mathbb{R}^{n}}|\nabla_{x}v(t,x)|^{2}dx\leq C(1-t)^{-1}$
and the virial estimate;
(2.19) $\int_{\mathbb{R}^{n}}|x|^{2}|v(t,x)|^{2}dx\leq C(1-t).$
A natural interpretation of this estimate is that the solution is spreading out (dispersing) with a very strong rate.
The crucial estimate (2.19) is obtained via appropriate half - pseudoconformal transformation
(2.20) $v(t,x)=e^{-\frac{ix^{2}}{4(1-t)}}(1-t)^{-n/4}Y\left(t,\frac{x}{\sqrt{1-t}}\right).$
Then $Y(t,y)$ satisfies a non - linear Schrödinger equation involving inverted harmonic oscillator. We use a combination between virial identity for this equation and energy estimate involving the vector field
(2.21) $\partial_{y}-i\frac{y}{2}$
that allows us to obtain
(2.22) $\int_{\mathbb{R}^{n}}|y|^{2}|Y(t,y)|^{2}dy\leq C$
and then we arrive at (2.19).
@@
For any $u_{0}\in\mathcal{F}(H^{1})$ there exists a unique $u_{+}\in L^{2},$ so that
@@
For any $u_{0}\in\mathcal{F}(H^{1})$ there exists a unique $u_{+}\in L^{2},$ so that
@@ 6.3. Conservation of the pseudoconformal energy
(6.21) $\frac{1}{4(1+t)}\frac{d}{dt}\|J(1+t)u(t)\|_{L^{2}}^{2}+\frac{d}{dt}\frac{n}{2(n+1)}\|u(t)\|_{L^{2(n+1)/n}}^{2(n+1)/n}=0$
@@
(6.22) $\frac{d}{dt}\left(\frac{1}{4(1+t)}\|J(1+t)u(t)\|_{L^{2}}^{2}+\frac{n}{2(n+1)}\|u(t)\|_{L^{2(n+1)/n}}^{2(n+1)/n}\right)$
@@ 8. Weak convergence of modified wave operators
Then we have (4.10) and we can refer to Corolary 4.1 that implies
@@
Then we have (4.10) and we can refer to Corolary 4.1 that implies

### v2

@@ Abstract.
The solution to this problem can be reformulated in terms of the completeness of wave operators. For $n=1$, we obtain affirmative answer, provided appropriate control on $L^{\infty}$ norm even for large initial data. For $n=2$ completeness is established under suitable control of the $L^{\infty}$ norm of the solution $u(t).$
In the case where no control on the $L^{\infty}$ norm is available, we establish similar asymptotic behavior by replacing the convergence in $L^{2}$ norm in (0.1) with a weaker convergence.
@@ 1.2. Purpose of the work
Our goal is to discuss Q2) without smallness on the initial data for $n=1,2$ or $n\geq 3$ and assuming $L^{\infty}$ control of the solution.
@@
Let us mention that in the case $n=1$ the question Q1) is treated in a preprint [22].
@@ 2. Main results
In the case $n=1$ we can control the $L^{\infty}$ norm and this is our next result.
@@
For any $R>0$ there exists $C(R)$ so that for any initial data satisfying
(2.12) $u_{0}\in\mathcal{H}^{1,1},\ \ \|u_{0}\|_{\mathcal{H}^{1,1}}\leq R$
the Cauchy problem
(2.13) $(i\partial_{t}+\frac{1}{2}\Delta_{x})u=u|u|^{2},\ \ t\geq 0;$
$u(0)=u_{0}\in\mathcal{H}^{1,1},$
has a unique global solution $u\in C([0,+\infty);\mathcal{H}^{1,1})$ such that
(2.14) $\|u(t,\cdot)\|_{L^{\infty}}\leq C(R)\ t^{-1/3},\ \ t\geq 1.$
In the case $n=2$ we can control the $L^{\infty}$ norm under the assumption that initial data are radial.
@@
For $n=2$ any $R>0$ there exists $C(R)$ so that for any initial data satisfying
(2.15) $u_{0}\in\mathcal{H}^{1,1}_{rad},\ \|u_{0}\|_{\mathcal{H}^{1,1}_{rad}}\leq R$
the Cauchy problem
(2.16) $(i\partial_{t}+\frac{1}{2}\Delta_{x})u=u|u|,\ \ t\geq 0;$
$u(0)=u_{0}\in\mathcal{H}^{1,1}_{rad},$
has a unique global solution $u\in C([0,+\infty);\mathcal{H}^{1,1}_{rad})$ such that
(2.17) $|u(t,x)|\leq C(R)\varepsilon^{-1/2}|x|^{-\varepsilon}t^{-(1-\varepsilon)/2}$
for any $x\in\mathbb{R}^{2}\setminus\{0\},$ any $t>1$ and for any $\varepsilon\in(0,1/2].$
The idea of the proofs of these $L^{\infty}$ bounds is based on the pseudo - conformal estimate
(2.18) $t^{-1}\|J(t)u(t)\|_{L^{2}}^{2}\lesssim 1$
where $t\geq 1$ and
(2.19) $J(t)=x+it\nabla=U(t)xU(-t)=M(t)it\nabla M(-t).$
@@
Let $n=2.$ For any $u_{0}\in\mathcal{F}(H^{1})$ there exists a unique $u_{+}\in L^{2},$ so that
@@
Let $n\geq 3.$ For any $u_{0}\in\mathcal{F}(H^{1})$ there exists a unique $u_{+}\in L^{2},$ so that
@@ 6.3. Conservation of the pseudoconformal energy
(6.21) $\frac{(1+t)^{2}}{4}\frac{d}{dt}\left(\frac{1}{1+t}\|J(1+t)u(t)\|_{L^{2}}^{2}\right)$
$+(1+t)^{2}\frac{d}{dt}\left(\frac{n(1+t)}{2(n+1)}\|u(t)\|_{L^{2(n+1)/n}}^{2(n+1)/n}\right)$
$+\frac{1}{4}\|J(1+t)u(t)\|_{L^{2}}^{2}=0$
@@
(6.22) $\frac{d}{dt}\left(\frac{1}{4(1+t)}\|J(1+t)u(t)\|_{L^{2}}^{2}+\frac{n(1+t)}{2(n+1)}\|u(t)\|_{L^{2(n+1)/n}}^{2(n+1)/n}\right)$
@@
8. $L^{\infty}$ bounds
We assume that $u$ is solution to the Cauchy problem (1.1) with initial data in $\mathcal{H}^{1,1}.$
We have the following conservation laws.
  • Charge:
    $\|u(t)\|_{L^{2}}^{2}=\|u_{0}\|_{L^{2}}^{2}.$
  • Energy :
    $\|\nabla u(t)\|_{L^{2}}^{2}+\frac{2n}{n+1}\|u(t)\|_{L^{2+2/n}}^{2+2/n}=\|\nabla u_{0}\|_{L^{2}}^{2}+\frac{2n}{n+1}\|u_{0}\|_{L^{2+2/n}}^{2+2/n}.$
  • Pseudo - conformal charge:
    $+\int_{0}^{t}(1+s)^{-2}\|J(s+1)u(s)\|_{L^{2}}^{2}ds$
    $=\|J(1)u(0)\|_{L^{2}}^{2}+\frac{2n}{n+1}\|u(0)\|_{L^{2+2/n}}^{2+2/n},$
    where
    (8.2) $J(t)=x+it\nabla=U(t)xU(-t)=M(t)it\nabla M(-t).$
Proof of Theorem 2.3.
Now we use the estimates  [^1] We use the following estimates. For any $\varphi\in H^{1}(\mathbb{R})$ (8.3) $|\varphi(x)|^{p}=-\int_{x}^{+\infty}\frac{d}{dy}\left(|\varphi(y)|^{p}\right)dy$ $\leq p\int_{-\infty}^{+\infty}|\varphi|^{p-1}|\varphi^{\prime}|dy\leq p\|\varphi\|_{L^{2p-2}}^{p-1}\|\varphi^{\prime}\|_{L^{2}}.\ \$Hence (8.4) $\|\varphi\|_{L^{\infty}}\leq p^{1/p}\|\varphi\|_{L^{2p-2}}^{1-1/p}\|\varphi^{\prime}\|_{L^{2}}^{1/p}.$
(8.5) $\|u(t)\|_{L^{\infty}}^{3}=\|M(-t)u(t)\|_{L^{\infty}}^{3}\leq 3\|M(-t)u(t)\|^{2}_{L^{4}}\|\nabla M(-t)u\|_{L^{2}}$
$=3\|u(t)\|^{2}_{L^{4}}\left(t^{-1}\|J(t)u(t)\|_{L^{2}}\right)\leq C(t^{-1/4})^{2}(t^{-1}t^{1/2})=Ct^{-1}$
for $t\geq 1.$ Hence we have (2.14). ∎
Proof of Theorem 2.4.
We take $u_{0}\in\mathcal{H}^{1,1}$ radial. Then for any $x\in\mathbb{R}^{2}\setminus\{0\}$, $p\geq 2,$ and $t\geq 1$ we have  [^2] Here for any $\varphi(r)\in H^{1}(\mathbb{R}),$ any $r>0$ and any $p\geq 2$ we use the inequalities (8.6) $|\varphi(r)|^{p}\leq p\left(\int_{r}^{+\infty}|\varphi(p)|^{2p-2}dp\right)^{1/2}\left(\int_{r}^{+\infty}|\varphi^{\prime}(p)|^{2}dp\right)^{1/2}$ $\leq\frac{p}{r}\left(\int_{r}^{+\infty}|\varphi(p)|^{2p-2}pdp\right)^{1/2}\left(\int_{r}^{+\infty}|\varphi^{\prime}(p)|^{2}pdp\right)^{1/2}.$ Hence, for $\varphi\in H^{1}_{rad}(\mathbb{R}^{2})$ and any $x\in\mathbb{R}^{2}\setminus\{0\}$ we have (8.7) $|\varphi(x)|^{p}\leq\frac{p}{|x|}\|\varphi\|_{L^{2p-2}}^{p-1}\|\nabla\varphi\|_{L^{2}}.$ On the other hand, there exists a constant $C\geq 1,$ so that for any $q\in[2,+\infty)$ and any $\varphi\in H^{1}(\mathbb{R}^{2})$ we have (8.8) $\|\varphi\|_{L^{q}}\leq Cq^{1/2}\|\varphi\|_{L^{2}}^{2/q}\|\nabla\varphi\|_{L^{2}}^{1-2/q}.$
(8.9) $|u(t,x)|^{p}=|M(-t)u(t,x)|^{p}\leq\frac{p}{|x|}\|M(-t)u(t)\|_{L^{2p-2}}^{p-1}\|\nabla M(-t)u(t)\|_{L^{2}}.$
Using the fact that
(8.10) $\|M(-t)u(t)\|_{L^{2p-2}}$
$\leq C(2p-2)^{1/2}\|M(-t)u(t)\|_{L^{2}}^{2/(2p-2)}\|\nabla M(-t)u(t)\|_{L^{2}}^{1-2/(2p-2)},$
we obtain
(8.11) $|u(t,x)|^{p}$
$\leq\frac{p}{|x|}C^{p-1}2^{(p-1)/2}p^{(p-1)/2}\|u_{0}\|_{L^{2}}\|\nabla M(-t)u(t)\|_{L^{2}}^{p-1}$
$\leq\frac{p}{|x|}C^{p-1}2^{(p-1)/2}p^{(p-1)/2}\|u_{0}\|_{L^{2}}\left(t^{-1}\|J(t)u(t)\|_{L^{2}}\right)^{p-1}$
$\leq\frac{p}{|x|t^{p-1}}C^{p}(2p)^{p/2}\|u_{0}\|_{L^{2}}\|J(t)u(t)\|_{L^{2}}^{p-1}.$
for any $t\geq 1.$ In this way, we find
(8.12) $|u(t,x)|\leq\left(\frac{p}{|x|}\|u_{0}\|_{L^{2}}\right)^{1/p}C(2p)^{1/2}t^{-(p-1)/(2p)}.$
Substitution $\varepsilon=1/p$ implies (2.17).
@@ 9. Weak convergence of modified wave operators
Then we have (4.10) and we can refer to Corollary 4.1 that implies
@@
Then we have (4.10) and we can refer to Corollary 4.1 that implies